OpenStudy (anonymous):

A Group of students are going on field trip. The transportation will include vans, which can hold 7 students each, and buses which can hold 25 students each, the school has five more vans than buses. If 127 students and 4 teachers are going on the trip, how many vans and buses will be used? steps, equation, anything to help.

OpenStudy (ash2326):

@Myammecutie are you here?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

Let there be V no. of vans and b no. of buses. Do you have relation between no. of both?

OpenStudy (anonymous):

there is five more vans than buses

OpenStudy (ash2326):

so how B and V are related? Could you write an expression for that in terms of B and V ?

OpenStudy (anonymous):

i need an equation but its not coming to me

OpenStudy (ash2326):

5 more van than buses so \[V=B+5\] do you get this?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

Now how many people are there totally? students+teachers

OpenStudy (anonymous):

131

OpenStudy (ash2326):

It's given in the question that a bus can hold 25 students and a van can hold 7 students each, so could you write an equation? hint= no. of people in bus+ van =131

OpenStudy (anonymous):

sooo 131=V7+B5

OpenStudy (ash2326):

yes \[7V+5B=131\] and we had found an equation earlier \[V=5+B\] could you solve both of these equations to find V and B?

OpenStudy (anonymous):

no im confused

OpenStudy (ash2326):

which part you didn't get?

OpenStudy (anonymous):

the equations i dont understand how to solve for two variables

OpenStudy (ash2326):

we have one equation V=B+5 and another 7V+25B=131 put V from first into second \[7(B+5)+25B=131\] \[7B+35+25B=131\] Can you find B from here?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i got 3

OpenStudy (anonymous):

buses

OpenStudy (ash2326):

right:D now could you find V?

OpenStudy (anonymous):

its 8

OpenStudy (ash2326):

Good:D so did you understand?

OpenStudy (anonymous):

Yes :)) thank you

OpenStudy (ash2326):

welcome:)